# Will Neural Scaling Laws Activate Jevons’ Paradox in AI Labor Markets? A Time-Varying Elasticity of Substitution (VES) Analysis

### v1

@@ 3 Exponential Declines in the AI Price, Scaling Laws, and Increased Elasticity of Substitution (VES)
Based on the model, as an aid to categorization, AI adoption could be thought of as occurring in five phases as presented below:
@@
    Phase 1 represents the complementarity-dominated regime where $\sigma(t)<1$ since $p_{A}<p_{H}=1$. AI enhances rather than replaces human production, with adoption driven primarily by initial price advantage. Specifically, when $t$ is near 0 market penetration is driven by both price effects and the initial low market share. In other words, AI firms must try to begin with a low price to initiate or jumpstart the market. During this phase, perhaps a long phase, AI goods, services, or tasks are complements to the human counterparts. Even if $p_{A}(t)$ is almost 0, human goods, services, or tasks will have positive share or $r_{A}<0.5$. Jevons’ Paradox does not occur during this phase. Currently, tasks such as coding display relatively large shares of AI, but under 50%. [^8] In the October 29th 2024 earnings call, Google CEO Sundar Pichai stated “We’re also using AI internally to improve our coding processes, which is boosting productivity and efficiency. Today, more than a quarter of all new code at Google is generated by AI, then reviewed and accepted by engineers.” In Patel (2025), Nadella states his beliefs in human-AI complementarity, but Nadella (2025) also promotes Jevons’ Paradox which indicates they are substitutes. Although both could be true, given factors outside of the simple model here via large-scale economic growth. Altman (2025) often suggests AI will exit this first phase.
@@
    Phase 2 marks the critical transition point where $\sigma(t)=1$, corresponding to unitary elasticity of substitution. This occurs when $t^{*}=\phi^{-1}$. This marks the tipping point where the elasticity of substitution between AI and human goods, services, or tasks is equal so that $\sigma(t)=1$. This phase exhibits properties analogous to Cobb-Douglas production, with constant quantity and revenue shares across AI and human inputs despite changing relative prices. In fact, this corresponds to $r_{A}=0.5$, a large market share. At this point, Jevons’ Paradox is at the cusp of occurring.
@@
    Phase 3 initiates true Jevons’ dynamics in revenue share where $\sigma(t)>1$ happens when $t>t^{*}$). The price elasticity satisfies $E_{AA}<-1$, indicating that percentage increases in quantity exceed percentage decreases in price. Jevons’ Paradox comes into play and price decreases actually increase revenue by spurring more usage. As $t$ increases, the term $-d\cdot\phi\cdot t^{2}$ dominates, representing the increasing influence of computational scaling and lower prices on market transformation. In this phase the AI market share is between 50 and 100% ($10.5<r_{A}<$ 1).
@@
    Phase 4 begins when $\sigma(t)\geq 2$ at $t\geq 2\cdot t^{*}$ and thus accentuates Jevons’ dynamics in revenue share. In the beginning of this phase, when $\sigma=2$, the revenue share shows a simple relation in (14) with $p_{A}$, the price of AI, based on substitution of $\sigma=2$ into (3). Note, if $p_{A}$ is already low by the time $\sigma=2$, the revenue share $r_{A}$ may already risen close to 1. The converse of (14) appears in (15) where the price $p_{A}$ shows a simple relation with revenue share, $r_{A}$. Also, the elasticity of revenue share to price $E_{r_{A},p_{A}}$ has a simple form in (16), based on substitution of $\sigma=2$ into (5). For low values of $p_{A}$, the elasticity may have already have low magnitude as lowering the price more will not bring forth more revenue.
    $r_{A}$ $=\dfrac{1}{1+p_{A}}$ (14)
    $p_{A}$ $=\frac{1-r_{A}}{r_{A}}$ (15)
    $E_{r_{A},p_{A}}$ $=-\dfrac{p_{A}}{{1+p_{A}}}=-p_{A}\cdot r_{A}$ (16)
@@
    Phase 5 represents market saturation within the constrained partial equilibrium framework, where $r_{A}=1$. At this boundary, Jevons’ Paradox no longer applies as the price elasticity equals $-1$, indicating unitary elasticity. Of course, if AI spurs economic growth the overall market could expand and Jevons’ Paradox could continue to hold. This sentiment has been put forth by industry leaders, but addressing this lies outside the scope of this simple model.
@@ 4 Conclusion
This analysis provides a framework that incorporates these statements to make the underlying logic clearer. Specifically, we examine market transitions to AI-produced goods using an elasticity of substitution framework, drawing on the identity presented in Gørtz (1977). We introduce a variable elasticity of substitution (VES) model where the increased performance of AI based on the scaling laws expands as a function of computational growth that leads to the elasticity of substitution $\sigma(t)$ growing over time. The growth in $\sigma(t)$ coupled with falling prices for AI driven by competition and the lower marginal costs of computation (stemming from hardware, algorithmic, and software progress) create an environment where AI becomes better and more inexpensive relative to humans over time.
@@
Having presented this model to understand the perspective of AI proponents, many alternative perspectives exist as to the reasons why explosive AI growth may or may not occur (Erdil and Besiroglu, 2024). The simple partial equilibrium model presented above does not account for the many positive or negative feedback loops in a more general equilibrium that could accelerate or impede adoption. For example, we held the price of the human product at 1, regardless of the context, whereas this may change in response to AI competition. We also do not discuss the differential changes in substitution and associated feedback loops that could occur across products, services, or tasks as AI, through the mechanism of high-quality data availability (e.g., programming, science). In addition, the optimistic view of AI stresses the increase in overall wealth and quality-of-life potential which could feedback into choices. Finally, if AI grows income and wealth, usage could continue to grow despite having captured the entire market which could lead to a continuation of Jevons’ Paradox. In addition, we do not discuss possible impediments to AI development such as declining quantity and quality of data, energy usage, capital costs, scaling laws failing to hold, frictions, and so forth.

### v2

@@ 3.2 Five Phase Categorization
Based on the model and prior discussion, as an aid to categorization, AI adoption could be thought of as occurring in five phases as presented below:
@@
[…]
@@
    Phase 2 marks the critical transition point where $\sigma(t)=1$, corresponding to unitary elasticity of substitution. This occurs when $t^{*}=\phi^{-1}$. This marks the tipping point where the elasticity of substitution between AI and human goods, services, or tasks is equal so that $\sigma(t)=1$. This phase exhibits properties analogous to Cobb-Douglas production, with constant expenditure shares across AI and human inputs despite changing relative prices. In fact, this corresponds to $r_{A}=\alpha$, which yet may be a small market share. Note, this occurs regardless of the price trajectory. At this point, Jevons’ Paradox is at the cusp of occurring.
@@
    Phase 3 initiates true Jevons’ dynamics in revenue share where $\sigma(t)>1$ happens when $t>t^{*}$. Jevons’ Paradox comes into play and price decreases actually increase revenue by spurring more usage. As $t$ increases, the term $-d\cdot\phi\cdot t^{2}$ dominates, representing the increasing influence of computational scaling and lower prices on market transformation. In this phase the AI market share is between $\alpha$ and 1.
@@
    Phase 4 begins when $\sigma(t)\geq 2$ at $t\geq 2\cdot t^{*}$ and thus accentuates Jevons’ dynamics in revenue share. In the beginning of this phase, when $\sigma=2$, the revenue share shows a simple relation in (27) with $p_{A}$, the price of AI, based on substitution of $\sigma=2$ into (2). Note, if $p_{A}$ is already low by the time $\sigma=2$, the revenue share $r_{A}$ may have already risen close to 1. The converse of (27) appears in (28) where the price $p_{A}$ shows a simple relation with revenue share, $r_{A}$. Also, the elasticity of revenue share to price $E_{r_{A},p_{A}}$ has a simple form in (29), based on substitution of $\sigma=2$ into (3). For low values of $p_{A}$, the elasticity may already have low magnitude as lowering the price more will not bring forth more revenue.
    $r_{A}$ $=\frac{\alpha}{(1-\alpha)p_{A}+\alpha}$ (27)
    $p_{A}$ $=\left(\frac{\alpha}{1-\alpha}\right)\cdot\left(\frac{1-r_{A}}{r_{A}}\right)$ (28)
    $E_{r_{A},p_{A}}$ $=\frac{\partial\ln r_{A}}{\partial\ln p_{A}}=-\bigl{(}\sigma-1\bigr{)}(1-r_{A})\Big{|}_{\sigma=2}\;=\;-\bigl{(}1-r_{A}\bigr{)}$ (29)
@@
    Phase 5 represents market saturation within the constrained partial equilibrium framework, where $r_{A}=1$. At this boundary, Jevons’ Paradox (as applied in terms of revenue share) no longer applies as the elasticity of revenue share goes to 0. Of course, if AI spurs economic growth the overall market could expand and lead to actual revenue to continue to rise. This sentiment has been put forth by industry leaders, but addressing this lies outside the scope of this simple model.
4 Illustration
To make the preceding analysis more immediate, we provide Figure 1 which shows the trajectory of various levels of AI revenue share as a function of the initial preference weight $\alpha=0.001$, initial price $p_{A}(0)=0.5$, and five visually distinct combinations of computation growth parameter $g$, price decline parameter $d$, and AI learning rate $\delta$ over a 30 year period. We selected the initial preference weight $\alpha=0.001$ for two reasons. First, this allows showing a transition or adoption curve going over the full range from $r_{A}(t)$ close to 0 for low $t$ and close to 1 (in some cases) for large $t$. Second, AI has always had a low pecuniary price, even at low adoption rates. The unusual nature of AI means that for an equal pecuniary price, it would have low rates of adoption. This could change over time with familiarity.
AI Revenue Share Dynamics with Time-Varying Elasticity of Substitution Refer to caption
Figure 1: Revenue-share paths generated from (9) with $\alpha=0.001$ and $p_{A}(0)=0.5$. After pruning visually indistinguishable cases, the surviving parameter sets are $(g,d,\delta)\in\{(0.5,0.5,0.05),\,(0.5,0.5,0.15),\,(0.5,1.5,0.15),\,(1.5,0.5,0.15),\,(1.5,1.5,0.15)\}$. Hollow markers depend only on the elasticity path $\sigma(t)=\phi\cdot t$ with $\phi=\delta\cdot g$: a circle at $t^{\ast}=1/\phi$ marks $\sigma=1$ (entry into the Jevons regime), while a cross at $2\cdot t^{\ast}$ marks $\sigma=2$ (strong Jevons). The price-decay rate $d$ shifts entire curves without moving these phase boundaries, separating the quality-driven phase structure from the price-driven climb in $r_{A}(t)$.
Before the circle marker the system is in Phase 1, where AI and human inputs are complements and Jevons’ Paradox does not arise. Between the circle and the cross ($21<\sigma<$ 2) the system enters Phase 3, exhibiting elastic demand and Jevons dynamics. To the right of the cross ($\sigma>2$) it moves into Phase 4, a “strong” Jevons regime. Phase 2 is a knife-edge at $\sigma=1$ and Phase 5 is the limiting saturation as $r_{A}(t)\to 1$. Thus the figure illustrates how identical substitution paths can yield different revenue trajectories when price declines occur at different speeds.
The two black markers depend only on the compound rate $\phi=\delta\cdot g$. Doubling $g$ (faster compute growth) or doubling $\delta$ (a steeper scaling law) halves the time needed to reach each phase boundary. Hence curves with the same $\phi$ share the same phase times regardless of how the product is split between the two components. By contrast, the price-decay rate $d$ merely translates the entire $r_{A}(t)$ path left or right without changing $t^{\ast}$ or $2\cdot t^{\ast}$. When $\phi$ is large even a modest $d$ yields a near-vertical takeover; when $\phi$ is small, aggressive price cuts (large $d$) are required just to move off the horizontal axis. In short, $\phi$ sets the clock for substitution, whereas $d$ determines how quickly revenue follows once substitution becomes technologically feasible.
The figure thus separates the quality-driven phase structure (dictated by $\sigma(t)$) from the price-driven rise in $r_{A}(t)$, illustrating how rapid cost declines ($d$ large) can accelerate revenue capture even when the elasticity path $\sigma(t)$ is common across scenarios.
Because such interest exists on AI timelines, we examine all the parameter combinations in Table 1. It displays the compound rate $\phi=\delta\cdot g$, the phase 2 boundary $t^{\ast}=1/\phi$, and the AI revenue share at $t=10$ and $t=30$ years. It reveals cases where AI never dominates (rows 1,3 ) and those where it does in as little as 10 years (last row).
$g$ $d$ $\delta$ $\phi$ $t^{\ast}$ $r_{A}(10)$ $r_{A}(30)$
$0.5$ $0.5$ $0.05$ $0.025$ $40.000$ 0.000 0140.0000140.000\,0140.000 014 0.000 0200.0000200.000\,0200.000 020
$0.5$ $0.5$ $0.15$ $0.075$ $13.333$ 0.000 2410.0002410.000\,2410.000 241 0.999 9970.9999970.999\,9970.999 997
$0.5$ $1.5$ $0.05$ $0.025$ $40.000$ 0.000 $0$.0000000.000\,$0$.000 000 0.000 $00\,0000$
$0.5$ $1.5$ $0.15$ $0.075$ $13.333$ 0.000 0200.0000200.000\,0200.000 020 1.000 0001.0000001.000\,0001.000 000
$1.5$ $0.5$ $0.05$ $0.075$ $13.333$ 0.000 2410.0002410.000\,2410.000 241 0.999 9970.9999970.999\,9970.999 997
$1.5$ $0.5$ $0.15$ $0.225$ $4.444$ 0.552 2290.5522290.552\,2290.552 229 1.000 0001.0000001.000\,0001.000 000
$1.5$ $1.5$ $0.05$ $0.075$ $13.333$ 0.000 0200.0000200.000\,0200.000 020 1.000 0001.0000001.000\,0001.000 000
$1.5$ $1.5$ $0.15$ $0.225$ $4.444$ 0.999 9970.9999970.999\,9970.999 997 1.000 0001.0000001.000\,0001.000 000
Table 1: Phase-transition time $t^{\ast}$ and AI revenue shares for each parameter triple, computed with $\alpha=0.001$ and $p_{A}(0)=0.5$.
We present some of the key insights below:
  1. 1.
    Phase timing depends on $\phi=\delta\cdot g$. Rows with the same $\phi$ share the same $t^{\ast}$, regardless of the price-decay rate $d$.
  2. 2.
    Price cuts affect revenue primarily after $\sigma>1$. Compare rows with identical $\phi$ but different $d$, such as row two and row four. The identical $\phi=0.075$ gives the same $t^{\ast}$. However, the effectiveness of aggressive price cuts (a higher $d$) in accelerating revenue share growth is primarily realized after $\sigma>1$. Before this threshold (i.e., when $t<t^{*}$), a higher $d$ can, under certain conditions illustrated here, lead to a temporarily lower revenue share $r_{A}(t)$ compared to a scenario with slower price declines.
  3. 3.
    Large $\phi$ ensures rapid saturation. When $\phi=0.225$ the revenue share already exceeds $55\%$ by year 10 and essentially saturates by year 30, independent of $d$.
  4. 4.
    Small $\phi$ plus aggressive price cuts is ineffective. The combination $\phi=0.025$, $d=1.5$ leaves $r_{A}(30)\approx 0$; Jevons’ Paradox never materializes without faster compute growth or a steeper scaling law.
Examination of the Figure and Table reveal that investment that raises either the compute growth rate $g$ or the scaling-law sensitivity $\delta$ offers far greater leverage over AI market penetration than slashing prices $d$ alone. Price competition matters chiefly after technology makes AI and human inputs substitutable. In other words, AI quality is the most important factor in the AI transition.
@@ 5 Conclusion
This analysis provides a framework that incorporates these statements to make the underlying logic clearer. Specifically, we examine market transitions to AI-produced goods using an elasticity of substitution framework. We introduce a variable elasticity of substitution (VES) model where the increased performance of AI based on the scaling laws expands as a function of computational growth that leads to the elasticity of substitution $\sigma(t)$ growing over time. The growth in $\sigma(t)$ coupled with falling prices for AI driven by competition and the lower marginal costs of computation (stemming from hardware, algorithmic, and software progress) create an environment where AI becomes better and more inexpensive relative to humans over time. This can still remain true under less restrictive assumptions, but these may lengthen the timelines.
@@
In Section 4 we provided some specific scenarios to make the analysis more immediate and illustrated these via Figure 1 and Table 1. These scenarios underscore a simple hierarchy. First, the compound rate $\phi=\delta\cdot g$ dictates when the economy reaches the substitution thresholds $\sigma=1$ and $\sigma=2$; doubling either the compute-growth rate $g$ or the scaling-law coefficient $\delta$ halves the time to each phase boundary. Second, once $\sigma>1$, the price decay rate $d$ controls how fast revenue flows to AI providers along an otherwise fixed substitution path. Finally, no amount of price cutting can compensate for a low $\phi$: when $\delta\cdot g$ is too small the Jevons regime may not arrive within any relevant planning horizon. For policymakers and investors this means that subsidizing algorithmic or hardware efficiency is a more potent lever than temporary price support, while for firms it suggests that competition on quality dominates price competition until AI and human inputs become bona fide substitutes.
However, since the translation of the gain from extra computation has a different effect by task, product, or service, the curves or scenarios depicted may also represent different share trajectories across individual tasks, products, or services. Insofar, as some scenarios never show an AI takeoff (perhaps martial art matches) while others show quick domination (perhaps coding), this indicates some of the disruption associated with AI will stem from optimization across continual shifting prices and capabilities.
For any given scenario, the steeper ‘S’-curve relative to a standard logistic, driven by the product $c$ of rates of price declines, computational growth, and quality improvements could further exacerbate the disruption making it seem abrupt or unanticipated.
@@
Having presented this model to understand the perspective of AI proponents, many alternative perspectives exist on the long term growth of AI (Erdil and Besiroglu, 2024). The simple partial equilibrium model presented above does not account for the many positive or negative feedback loops in a more general equilibrium that could accelerate or impede adoption. For example, we held the price of the human product at 1, regardless of the context, whereas this may change in response to AI competition. We also provide a perhaps false choice as AI versus human when hybridization has obvious potential. We also do not discuss the differential changes in substitution and associated feedback loops that could occur across products, services, or tasks as AI, through the mechanism of high-quality data availability (e.g., programming, science). Similarly, we neglect network and lock-in effects which could vary somewhat the trajectory. We treat quality gains as shifts in $\sigma(t)$ and cost declines as $p_{A}(t)$. Yet scaling laws plausibly do both together (computational gains drive both). In addition, the optimistic view of AI stresses the increase in overall wealth and quality-of-life potential which could feedback into choices. If AI grows income and wealth, usage could continue to grow despite having captured the entire market which could lead to a continuation of Jevons’ Paradox. In addition, we do not discuss possible impediments to AI development such as declining quantity and quality of data, energy usage, capital costs, scaling laws failing to hold, frictions, regulation, and other possible bottlenecks. Early progress could be misleading as AI can initially substitute in easy settings, but may encounter a large range of more difficult settings later.
The education sector offers an interesting counterpoint to tech industry invocations of Jevons’ Paradox. A number of public and accredited open universities exist across the world that allow credit by examination and other alternative ways of measuring knowledge acquisition such as Thomas Edison State University (US), the Open University (UK), Indira Gandhi National Open University (IN), and others. These institutions have delivered education at fraction of traditional higher education costs for decades. The internet reduced content distribution costs to essentially zero, yet market transformation never materialized. TESU captures negligible US market share with under 12,000 students and the even more successful Open University has captured only a plurality of the UK market. Under the proposed framework, this reveals education as a $\sigma<1$ market where Jevons’ Paradox fails to manifest in revenue terms. AI advocates suggest their technology transcends mere pecuniary cost reduction, offering superior knowledge, tireless availability, and adaptive personalization. However, the current analysis indicates that these quality improvements must push the elasticity of substitution above unity before capturing significant revenue share. The success of AI-driven platforms like Duolingo in specific educational niches suggests this threshold is achievable, but the resilience of traditional universities despite enormous cost disadvantages implies that broad educational transformation requires AI to deliver substitutability improvements exceeding current capabilities. Undoubtedly, part of the current lack of substitution comes from the different experience of being around other people of the same age in similar circumstances that help create substantial network effects for graduates as well as other non-pecuniary aspects of university life. This current Maginot Line of higher education may yet fall as AI finds ways to circumvent it, much as social media created new forms of connection rather than replicating traditional social experiences.